show thats the maximum magnitude e max of the elctric field along the axis of a uniformly charged ring occurs at x=a/2 and the vakues of q divide by6
1
Expert's answer
2020-12-02T07:33:31-0500

Answer

Electric field on the axis of ring is given by

"E=\frac{kqx}{(x^2+R^2)^{1.5}}"

For maximum of electric field differentiate above electric field with respect to x

"\frac{dE}{dx}=0"

then we get

"x=\pm\frac{R}{\sqrt{2}}" Here electric field is maximum.

Putting value of x then we get

"E_{max}=\frac{kqx}{((\frac{R}{\sqrt{2}})^2+R^2)^{1.5}}=\frac{2kq}{3\sqrt{3}R^2}"